# Learning with Statistical Equality Constraints

### v1

@@ B.4 Dual estimation error
If we expand $\hat{L}(f_{\theta^{(t)}},\gamma^{(t)})$ and $\hat{L}(f_{\theta^{(t)}},\gamma^{\prime})$ in terms of the cost-constraint vector $\hat{F}$ and the dual variables $\gamma^{(t)},\gamma^{\prime}$, the terms correspond to the objective vanish, and the remainder is $\sum_{k=1}^{K}(\gamma^{(t)}_{k}-\gamma^{\prime}_{k})\hat{F}_{k}(\theta^{(t)})$. Bringing $\hat{q}(\gamma^{\prime})$ to the RHS completes the proof. ∎
@@
Define $\cdot]_{+}:\mathbb{R}^{m}\to\mathbb{R}_{\geq 0}^{m}$ such that $\mathbf{x}]_{+}=\max{}\left(\mathbf{x},\mathbf{0}\right)$, where the maximum is taken componentwise. Let $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, then it follows that,
@@
Recall from Lemma C.3 that the following holds for arbitrary $t$.
@@
(74) proves that $U_{t}<U_{t-1}$ since $2\eta\rho<0$.
@@
Note that the proof requires the extra margin $(\alpha-1)\rho$ to show descent. Once $\hat{q}(\gamma^{(t-1)})-\hat{D}^{\star}\leq\rho$, (73) cannot prove descent because both terms are non-negative. Note the dependence of $T_{\max}$ on $\eta$ and $\rho$. While interpreting $T_{0}$ and $T_{\max}$, it is important to note that we cannot detect $T_{0}$ since we have neither $\hat{D}^{\star}$ nor $\hat{q}$ (since our oracle is approximate). However, we can compute $T_{\max}$, which is simply an upper bound on $T_{0}$. Proposition C.4 does not tell us what happens between $T_{0}$ and $T_{\max}$, in particular it does not tell us if our iterates become worse, either with respect to the mean squared error or dual function. While giving a guarantee for the last iterate (at $T=T_{\max})$ is challenging, we can give a guarantee for the average of the Lagrangian iterates. The following can also be written as a randomized result, such as in [10]. See 4.1
@@
Now since $\eta\leq\frac{2\alpha_{\eta}\rho}{(I+J)B^{2}}\Rightarrow\eta(I+J)B^{2}\leq 2\alpha_{\eta}\rho$, and $T\geq\frac{U_{0}\beta_{T}}{2\eta\rho}\Rightarrow\frac{U_{0}}{2\eta T}\leq\frac{\rho}{\beta_{T}}$, therefore, (80) reduces to,
$\hat{D}^{\star}-\frac{1}{T}\sum_{t=0}^{T-1}\hat{L}(f_{\theta^{(t)}},\gamma^{(t)})\leq\rho(1+\frac{1}{\beta_{T}}+\alpha_{\eta}),$
@@ D.1 Example : minimum norm interpolation
The feasibility and optimality sets of (Pi) coincide with (Pe) when $\epsilon=0$. However, the dual problems of both problems are different. In particular, the dual solution of (Pi) diverges as $\epsilon\to 0$ and the dual problem of (Pi) becomes ill-posed. However, the dual problem for (Pe) does not depend on $\epsilon$, and is well posed. Note that this also implies that recovering a good approximation of the solution to (Pe) using an unconstrained objective $\left\|w\right\|_{2}^{2}+\lambda\left\|Xw-y\right\|_{2}^{2}$ would need a very large weight $\lambda$.
@@
Now, since $g(w)$ is convex, and clearly continuous, therefore its sub-level set $g(w)\leq\epsilon$ is closed [50, Theorem 7.1]. Clearly, the set $\left\{w\;|\;\frac{1}{2}\left\|w\right\|_{2}^{2}\leq P^{\star}_{i}+\delta\right\}$ is nothing but the norm ball $B(0,\sqrt{2P^{\star}_{i}+\delta})$, and is closed and compact. Therefore $\text{Feas}(\text{P}_{i2})$ is compact. Setting $\delta=\gamma$ in (81) proves that $\text{Feas}(\text{P}_{i2})$ is non-empty as well. Since the objective, $\frac{1}{2}\left\|w\right\|_{2}^{2}$ is continuous, the compact non-emptiness of the feasibility set implies that $\text{Opt}(\text{P}_{i2})\neq\emptyset$ (see [51, Theorem 4.1.6]).
@@ D.2 Comparing equalities with double sided approximation
The difference between ($\text{P}_{0}$) and ($\text{P}_{\epsilon}$) is in the constraints–($\text{P}_{\epsilon}$) has twice the number of constraints as ($\text{P}_{0}$), and they are all inequality constraints, as opposed to the equality constraints in ($\text{P}_{0}$). This implies that both problems have different dual variables. Let $\mu\in\mathbb{R}^{J}$ denote the dual variable for ($\text{P}_{0}$) and let $\mu_{+},\mu_{-}\in\mathbb{R}_{\geq 0}^{J}$ denote the dual variables for ($\text{P}_{\epsilon}$), with ++ corresponding to the upper bound and - corresponding to the lower bound.
@@ E.1 Fairness
The fairness experiments were run on CPU-only nodes (AMD EPYC 7713) with 32 threads, with about 10-20 runs in parallel. Each run took at most 20m (when the termination condition was not met).
@@ E.2 Boundary value problems
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded connected region with boundary $\partial\Omega$, and define the domain of a BVP as $\mathcal{D}=\Omega\times(0,T]$ (where $T\in\mathbb{R}_{\geq 0}$) and let $\mathcal{H}=\left\{f_{\theta}\;|\;\theta\in\Theta\right\}$ be a set of functions defined on $\mathcal{D}$. A BVP is typically posed as the following problem : We call the three constraints the partial differential equation (PDE), boundary condition (BC) and initial condition (IC) respectively. As discussed in Section 2, the PDE constraint can be transformed to the constraint $\mathbb{E}_{\mathbb{P}_{p}}[(D[f_{\theta}](\mathbf{x},t)-\uptau(\mathbf{x},t))^{2}]=0$ for some distribution $\mathbb{P}_{p}$ that has support over the entire domain $\mathcal{D}$. The difference between this and the original formulation is essentially that the latter is with respect to the $L^{2}$ norm, while the former is a pointwise constraint, or equivalently over the $L^{\infty}$ norm with respect to any distribution whose support contains $\mathcal{D}$. The BC and IC constraints can similarly be transformed into their counterparts in (P-BVP) by picking $\mathbb{P}_{b}$ and $\mathbb{P}_{i}$ as distributions over $\partial\Omega\times(0,T)$ and $\Omega$ respectively.
@@
The BVP experiments were run (one run at a time) on an accelerated node (AMD EPYC 7713, Nvidia A40 GPU) with 16 CPU threads and one GPU. Both PINN and (P-C) took approximately 1h40m for each run (300k epochs), with negligible overhead for the constrained method.
@@ E.3 Interpolating classifiers
The interpolation experiments were run (one run at a time) on a workstation (AMD Ryzen 5 7600X 6-Core Processor, NVIDIA GeForce RTX 5090) with 12 CPU threads and one GPU. On CIFAR-10, ERM took about 16m while (P-CI) took about 18m. On CIFAR-100, the overhead was larger, with ERM still taking about 16m while the constrained method took out 31m. This may be mitigated by a vectorised implementation of the Lagrangian, but is left for future work.

### v2

@@ B.4 Dual estimation error
If we expand $\hat{L}(f_{\theta^{(t)}},\gamma^{(t)})$ and $\hat{L}(f_{\theta^{(t)}},\gamma^{\prime})$, then the terms in the Lagrangians correspond to the objective cancel out. We are left with the product of the constraint functions ($\hat{C}_{1}(\theta^{(t)}),\ldots,\hat{C}_{I+J}(\theta^{(t)})$) and the difference in dual variables $\gamma^{(t)},\gamma^{\prime}$, i.e. $\sum_{k=1}^{I+J}(\gamma^{(t)}_{k}-\gamma^{\prime}_{k})\hat{C}_{k}(\theta^{(t)})$. Bringing $\hat{q}(\gamma^{\prime})$ to the RHS completes the proof. ∎
@@
Define $\cdot]_{+}:\mathbb{R}^{m}\to\mathbb{R}_{+}^{m}$ such that $\mathbf{x}]_{+}=\max{}\left(\mathbf{x},\mathbf{0}\right)$, where the maximum is taken componentwise. Let $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, then it follows that,
@@
Lemma C.3 proves the following for arbitrary $t$,
@@
Equation (74) proves that $U_{t}<U_{t-1}$ since $2\eta\rho<0$.
@@
Note that the proof requires that $\alpha>1$ to show descent. Once $\hat{q}(\gamma^{(t-1)})-\hat{D}^{\star}\leq\rho$, (73) cannot prove descent because both terms are non-negative. Note the dependence of $T_{\max}$ on $\eta$ and $\rho$. While interpreting $T_{0}$ and $T_{\max}$, it is important to note that we cannot detect $T_{0}$ since we have neither $\hat{D}^{\star}$ nor $\hat{q}$ (since our oracle is approximate). However, we can compute $T_{\max}$, which is simply an upper bound on $T_{0}$. Proposition C.4 does not tell us what happens between $T_{0}$ and $T_{\max}$, in particular it does not tell us if our iterates become worse, either with respect to the mean squared error or dual function. While giving a guarantee for the last iterate (at $T=T_{\max})$ is challenging, we can give a guarantee for the average of the Lagrangian iterates. The following can also be written as a randomized result, such as in [10]. See 4.1
@@
Now since $\eta\leq\frac{\rho}{(I+J)B^{2}}\Rightarrow\frac{1}{2}\eta(I+J)B^{2}\leq\frac{1}{2}\rho$, and $T\geq\frac{U_{0}}{\eta\rho}\Rightarrow\frac{U_{0}}{2\eta T}\leq\frac{\rho}{2}$, therefore, (80) reduces to,
$\hat{D}^{\star}-\frac{1}{T}\sum_{t=0}^{T-1}\hat{L}(f_{\theta^{(t)}},\gamma^{(t)})\leq 2\rho,$
@@ D.1 Example : minimum norm interpolation
The feasibility and optimality sets of (Pi) coincide with (Pe) when $\epsilon=0$. However, the dual problems of both problems are different. The dual problem for (Pe) does not depend on $\epsilon$, and is well posed. However, the dual solution of (Pi) diverges as $\epsilon\to 0$ and the dual problem of (Pi) becomes ill-posed. Note that this also implies that recovering a good approximation of the solution to (Pe) using an unconstrained objective $\left\|w\right\|_{2}^{2}+\lambda\left\|Xw-y\right\|_{2}^{2}$ would need a very large weight $\lambda$.
@@
Now, since $g(w)$ is convex, and clearly continuous, therefore its sub-level set $g(w)\leq\epsilon$ is closed [50, Theorem 7.1]. Clearly, the set $\left\{w\;|\;\frac{1}{2}\left\|w\right\|_{2}^{2}\leq P^{\star}_{i}+\delta\right\}$ is nothing but the norm ball $B(0,\sqrt{2(P^{\star}_{i}+\delta}))$, and is closed and compact. Therefore $\text{Feas}(\text{P}_{i2})$ is compact. Setting $\delta=\gamma$ in (81) proves that $\text{Feas}(\text{P}_{i2})$ is non-empty as well. Since the objective, $\frac{1}{2}\left\|w\right\|_{2}^{2}$ is continuous, the compact non-emptiness of the feasibility set implies that $\text{Opt}(\text{P}_{i2})\neq\emptyset$ (see [51, Theorem 4.1.6]).
@@ D.2 Comparing equalities with double sided approximation
The difference between ($\text{P}_{0}$) and ($\text{P}_{\epsilon}$) is in the constraints–($\text{P}_{\epsilon}$) has twice the number of constraints as ($\text{P}_{0}$), and they are all inequality constraints, as opposed to the equality constraints in ($\text{P}_{0}$). This implies that both problems have different dual variables. Let $\mu\in\mathbb{R}^{J}$ denote the dual variable for ($\text{P}_{0}$) and let $\mu_{+},\mu_{-}\in\mathbb{R}_{+}^{J}$ denote the dual variables for ($\text{P}_{\epsilon}$), with ++ corresponding to the upper bound and - corresponding to the lower bound.
@@ E.1 Fairness
The fairness experiments were run on CPU-only nodes with 32 threads, with about 10-20 runs in parallel. Each run took at most 20m (when the termination condition was not met).
@@ E.2 Boundary value problems
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded connected region with boundary $\partial\Omega$, and define the domain of a BVP as $\mathcal{D}=\Omega\times(0,T]$ (where $T\in\mathbb{R}_{+}$) and let $\mathcal{H}=\left\{f_{\theta}\;|\;\theta\in\Theta\right\}$ be a set of functions defined on $\mathcal{D}$. A BVP is typically posed as the following problem : We call the three constraints the partial differential equation (PDE), boundary condition (BC) and initial condition (IC) respectively. As discussed in Section 2, the PDE constraint can be transformed to the constraint $\mathbb{E}_{\mathbb{P}_{p}}[(D[f_{\theta}](\mathbf{x},t)-\uptau(\mathbf{x},t))^{2}]=0$ for some distribution $\mathbb{P}_{p}$ that has support over the entire domain $\mathcal{D}$. The difference between this and the original formulation is essentially that the latter is with respect to the $L^{2}$ norm, while the former is a pointwise constraint, or equivalently over the $L^{\infty}$ norm with respect to any distribution whose support contains $\mathcal{D}$. The BC and IC constraints can similarly be transformed into their counterparts in (P-BVP) by picking $\mathbb{P}_{b}$ and $\mathbb{P}_{i}$ as distributions over $\partial\Omega\times(0,T)$ and $\Omega$ respectively.
@@
The BVP experiments were run (one run at a time) on an accelerated node with 16 CPU threads and one GPU. Both PINN and (P-C) took approximately 1h40m for each run (300k epochs), with negligible overhead for the constrained method.
@@ E.3 Interpolating classifiers
The interpolation experiments were run (one run at a time) on a workstation with 12 CPU threads and one GPU. On CIFAR-10, ERM took about 16m while (P-CI) took about 18m. On CIFAR-100, the overhead was larger, with ERM still taking about 16m while the constrained method took out 31m. This may be mitigated by a vectorised implementation of the Lagrangian, but is left for future work.